# UnlearnCanvas: A Stylized Image Dataset to Benchmark Machine Unlearning for Diffusion Models

### v2

@@ 5 Evaluation Pipeline
Phase II: Machine unlearning (Fig. 5b). Next, we utilize the selected MU methods for benchmarking to update the DM acquired in Phase I, aiming to eliminate a designated unlearning target (e.g., ‘Van Gogh style’). This results in the unlearned model (denoted as ${\bm{\theta}}_{\mathrm{u}}$), which is the DM updated to avoid generating the unlearning target. This unlearned DM will be used for unlearning performance evaluation.
Phase III: Answer set generation (Fig. 5c). Further, we utilize the unlearned model ${\bm{\theta}}_{\mathrm{u}}$ to generate a set of images conditioned on both the unlearning-related prompts and other innocent prompts. To comprehensively evaluate unlearning performance, three types of answer sets are generated:
Unlearning effectiveness: Images are generated in response to prompts that involve the unlearning target (e.g., any prompt like ‘An image in Van Gogh style’ for the case in Fig. 5). A successfully unlearned model will avoid generating images that match the unlearning target.
In-domain retainability: Images are generated with prompts in the same ‘domain’ as the unlearning target (e.g., style-related prompts for style unlearning), where the prompts can cover all the other styles provided in UnlearnCanvas except the one to be unlearned, e.g., ‘Van Gogh style’ in Fig. 5. This assesses the unlearned model’s ability to retain generation quality within the same domain.
Cross-domain retainability: Images are generated with prompts in other domains. For instance, in the case of style unlearning as shown in Fig. 5, the object domain is considered separate from the style domain. Prompts related to objects are used to evaluate the DM’s cross-domain generation capability after style unlearning.
@@
Table 3: Performance overview of different MU methods evaluated with UnlearnCanvas dataset. The performance assessment includes unlearning accuracy (UA), in-domain retain accuracy (IRA), cross-domain retain accuracy (CRA), and FID. Symbols $\uparrow$ or $\downarrow$ denote whether larger or smaller values represent better performance. Results are averaged over all the style and object unlearning cases. The best performance regarding each metric is highlighted in bold.
@@
ESD (Gandikota et al., 2023a) $\mathbf{98.58}\%$ $80.97\%$ $93.96\%$ $92.15\%$ $55.78\%$ $44.23\%$ $65.55$ 6163 17.8 4.3
FMN (Zhang et al., 2023a) $88.48\%$ $56.77\%$ $46.60\%$ $45.64\%$ $90.63\%$ $73.46\%$ $131.37$ $\mathbf{350}$ 17.9 4.2
UCE (Gandikota et al., 2023b) $98.40\%$ $60.22\%$ $47.71\%$ $\mathbf{94.31}\%$ $39.35\%$ $34.67\%$ $182.01$ 434 $\mathbf{5.1}$ $\mathbf{1.7}$
CA (Kumari et al., 2023) $60.82\%$ $\mathbf{96.01}\%$ $92.70\%$ $46.67\%$ $90.11\%$ $81.97\%$ $\mathbf{54.21}$ 734 10.1 4.2
SalUn (Fan et al., 2023) $86.26\%$ $90.39\%$ $\mathbf{95.08}\%$ $86.91\%$ $\mathbf{96.35}\%$ $\mathbf{99.59}\%$ $61.05$ 667 30.8 4.0
Phase IV: MU performance assessment (Fig. 5d). After Phase III, the answer set undergoes style/object classification for unlearning performance assessment. This classification results in three quantitative metrics ①-③.
Unlearning accuracy (UA): It represents the proportion of images generated by the unlearned DM using the unlearning target-related prompt (e.g., the Van Gogh style prompt in Fig. 5), which are not correctly classified into the corresponding class for the case. Therefore, a higher UA indicates better unlearning performance in preventing image generation from the unlearning target-related prompts.
In-domain retain accuracy (IRA): It is given by the classification accuracy of images generated by the unlearned DM using innocent prompts (not relevant to the unlearning target) in the same domain, e.g., all styles other than the Van Gogh style in Fig. 5.
Cross-domain retain accuracy (CRA): Similar to IRA, it is given by the classification accuracy of images generated by the unlearned DM using innocent prompts but in different domains, e.g., all object categories in Fig. 5.
In addition to the above accuracy metrics, we also utilize the FID score (④) to evaluate the distribution-wise generation quality of the unlearned DM. Furthermore, we will also monitor the efficiency of the unlearning process, taking into account factors such as run-time (⑤), storage space requirements (⑥), and memory costs (⑦). In total, the performance evaluation involves 7 metrics (①-⑦).
MU methods for DMs to be benchmarked. In this work, we assess the 5 most recently proposed MU methods for DMs, including ESD (Gandikota et al., 2023a), FMN (Zhang et al., 2023a), UCE (Gandikota et al., 2023b), CA (Kumari et al., 2023), and Salun (Fan et al., 2023). We remark that SA (Heng & Soh, 2023), a recently proposed method, is excluded from our evaluation due to its excessive computational resource requirements and time consumption. Detailed training settings for each unlearning method can be found in Appx. A.
Figure 6: Performance visualization of Table 3. For UA, IRA, and CRA, the results are averaged over all the style and object unlearning scenarios. Other metrics undertake the inverse operation as a smaller values represent better performance. Results are normalized to $0\%\sim 100\%$ per metric.
Figure 7: Left: Heatmap visualization of the unlearning and retain accuracy of ESD on UnlearnCanvas. The $x$-axis shows the tested concepts for image generation using the unlearned model, while the $y$-axis indicates the unlearning target. Concept types are distinguished by color: styles in blue and objects in orange. The figure is separated into different regions to represent corresponding evaluation metrics and the unlearning scopes ($A$ for UA, $B$ for IRA, $C$ for CRA; ‘1’ for style unlearning, ‘2’ for object unlearning). Diagonal regions ($A1$ and $A2$) indicate unlearning accuracy, and off-diagonal values ($B1$, $B2$, $C1$, and $C2$) represent retain accuracy. Higher values in lighter color denote better performance. The first row serves as a reference for comparison before unlearning. Zooming into the figure is recommended for detailed observation. Right: Representative cases illustrating each region with images generated before and after unlearning a specific concept.
6 Experiment Results
In this section, extensive experiments are conducted on UnlearnCanvas following the evaluation pipeline in Sec. 5. We show that previous MU evaluation, primarily focused on unlearning effectiveness, may present a biased view without considering retainability assessment (Tab. 3). We also find that preserving cross-domain retainability, as quantified by CRA, is more challenging than in-domain retainability (measured by IRA), a previously overlooked aspect in the literature. Furthermore, we show that no single MU method consistently excels across all metrics, emphasizing the potential for advancements in existing methods (Fig. 6). In addition, we present a detailed case study on ESD, showcasing how our benchmark offers a comprehensive analysis of an unlearning method’s mechanism on forgetting and retaining information (Fig. 7). Last but not least, we show that different MU methods display distinct unlearning characteristics, which are inherently linked to their underlying unlearning mechanisms (Fig. 8).
Overall performance assessment of MU for DMs. In Tab. 3, we provide an overview of the performance of existing unlearning methods, using the metrics discussed in Sec. 5. The key results are summarized below.
(1) Retainability is essential for comprehensive assessment of MU in DMs. Previous evaluations heavily relied on UA (unlearning accuracy), but our assessment reveals a crucial insight: relying solely on UA can provide a skewed view of MU method performance. For example, when unlearning artistic painting styles (that we call style unlearning), ESD and UCE achieve similar UA (both around $98\%$), but their retainability (IRA and CRA) differs significantly, with one over $80\%$ and the other around $60\%$. Similar conclusions apply to unlearning an image object (that we call object unlearning). These findings emphasize the importance of including retainability metrics in the evaluation framework, as demonstrated by UnlearnCanvas.
@@ 6 Experiment Results
(2) CRA is harder to retain than IRA. In both style and object unlearning, CRA presents a greater challenge for MU methods compared to IRA. This trend is consistent across various methods. For example, the UCE method (Gandikota et al., 2023b) exhibits a significant gap between its IRA and CRA, both in style ($60.22\%$ vs. $47.71\%$) and object ($39.35\%$ vs. $34.67\%$) unlearning scenarios. Similar patterns are observed with other methods like FMN and CA. This pattern suggests that while MU methods are somewhat effective at preserving concepts within the same domain (in terms of IRA), they face greater challenges in maintaining performance on unlearning target-unrelated concepts across domains (CRA), which has been overlooked in previous studies due to the absence of a multi-label dataset like UnlearnCanvas for systematic unlearning analysis.
(3) One MU method can perform differently in different domains. Take the FMN method (Gandikota et al., 2023b) as an example. It exhibits an evident performance difference between the two domains, achieving a UA of $88.48\%$ in style unlearning but a much lower UA of $45.64\%$ in object unlearning. In addition, its IRA and CRA metrics show a notable gap, with an IRA of $90.63\%$ for object unlearning compared to $56.77\%$ for style unlearning. These examples highlight that a method’s effectiveness in unlearning is not consistent across different scopes.
(4) No single method can excel in all aspects. Each method exhibits unique strengths and weaknesses. For example, while ESD (Gandikota et al., 2023a) typically leads an improvement in UA, it may lag behind in terms of IRA and CRA, e.g., in the case of object unlearning. Also, the SalUn method (Fan et al., 2023) shows exceptional retainability, with a CRA of $99.59\%$ in object unlearning, but its UA in the same category is lower at $86.91\%$. Furthermore, the UCE method (Kumari et al., 2023) enjoys remarkably high UA and efficiency but suffers from degradation in FID and retainability. For better visualization, in Fig. 6, we show this pattern by plotting the performance of different metrics reported in Tab. 1, illustrating that no method dominates across all radar chart dimensions. This underscores the complexity of unlearning in generative models and the need for further advancement.
@@
A closer look into retainability: A case study on ESD. We next perform a more in-depth analysis of DMs’ retainability after unlearning, with a focus on the ESD method (Gandikota et al., 2023a), which has shown superiority in Tab. 3 for style unlearning. In Fig. 7 (left), we present a heatmap that illustrates the per-style/object unlearning accuracy (diagonal) and retain accuracy (off-diagonal) for ESD. The heatmap segments are labeled to indicate different regions of interest (ROIs) related to our evaluation metrics: A1 and A2 for Unlearning Accuracy (UA) in style and object unlearning, B1 and B2 for In-Domain Retain Accuracy (IRA) in style and object unlearning, and C1 and C2 for Cross-Domain Retain Accuracy (CRA) in style and object unlearning. We highlight the following observations.
First, we observe distinct behaviors of ESD in unlearning styles compared to objects. The ROI C2 appears darker than C1, suggesting a reduced ability to retain styles during the unlearning of objects. Similar patterns are observable between B2 and B1. These findings indicate that an MU method might demonstrate a preference within a specific domain but face challenges in others. For a clearer explanation, Fig. 7 (right) provides image generation examples of the DM before and after unlearning within different ROIs.
Second, we observe that style/object unlearning is relatively easier compared to retaining the generation performance of unlearned DMs conditioned on unlearning-unrelated prompts. This is evident in the contrast between ROIs A1/A2 and ROIs B1/B2 or C1/C2. One exception is the case of unlearning the object ‘Sea’, which exhibits relatively low UA in A2. This may be attributed to the image generation post-unlearning still resembling sea waves, as shown by the image examples in Fig. 7 (right) for A2.
Third, we present a similar accuracy heatmap for SalUn in Fig. A1. In comparison to ESD, SalUn demonstrates more consistent performance across different unlearning scenarios. However, SalUn does not achieve the same level of UA as ESD, as evidenced by the lighter diagonal values in Fig. A1. This observation underscores the trade-off between unlearning effectiveness and retainability of MU in DMs.

### v4

@@ 3 Our Proposal: UnlearnCanvas Dataset
Phase II: Machine unlearning (Fig. 5 (b)). Next, we utilize a given MU method to update the DM acquired in Phase I, aiming to eliminate a designated unlearning target (e.g., Van Gogh style). This results in the unlearned DM (denoted as ${\bm{\theta}}_{\mathrm{u}}$), which avoids generating images in response to Van Gogh style-specific prompts.
Phase III: Answer set generation (Fig. 5 (c)). Further, we utilize the unlearned model ${\bm{\theta}}_{\mathrm{u}}$ to generate a set of images conditioned on both the unlearning-related prompts and other benign prompts. To comprehensively evaluate unlearning performance, three types of answer sets are generated. (a) To assess unlearning effectiveness, images are generated in response to prompts that involve the unlearning target (e.g., any prompt like ‘An image in Van Gogh style’ for the case in Fig. 5). (b) To assess in-domain retainability (i.e., the unlearned model’s ability to retain generation quality within the same domain), images are generated with prompts in the same domain as the unlearning target (e.g., style-related prompts for style unlearning). These prompts cover all other styles provided in UnlearnCanvas except the one to be unlearned. (c) To measure cross-domain retainability, images are generated with prompts in other domains. For instance, in the case of style unlearning as shown in Fig. 5, the object domain is considered separate from the style domain. Prompts related to objects are used to evaluate the DM’s cross-domain generation capability after style unlearning.
Phase IV: MU performance assessment (Fig. 5 (d)). After Phase III, the answer set undergoes style/object classification for unlearning performance assessment. This classification results in three quantitative metrics ①-③. ① Unlearning accuracy (UA): This represents the proportion of images generated by the unlearned DM using the unlearning target-related prompt that is not correctly classified into the corresponding class. A higher UA indicates better unlearning performance in preventing image generation from the unlearning target-related prompts. ② In-domain retain accuracy (IRA): This is the classification accuracy of images generated by the unlearned DM using innocent prompts (not relevant to the unlearning target) within the same domain. ③ Cross-domain retain accuracy (CRA): Similar to IRA, this is the classification accuracy of images generated by the unlearned DM using innocent prompts in different domains. In addition to the accuracy metrics, we also use the FID score (④) to evaluate the distribution-wise generation quality of the unlearned DM. Furthermore, we monitor the efficiency of the unlearning process, considering factors such as run-time (⑤), storage space requirements (⑥), and memory costs (⑦).
Figure 5: An illustration of the evaluation pipeline proposed in this work using UnlearnCanvas when unlearning a specific target concept ‘Van Gogh Style’. Unlearning performances (including the unlearning effectiveness and retainability) are quantitatively assessed (marked in blue) to accurately reflect the unlearning performance portrait. The unlearning target of the pipeline could traverse all the styles and objects to achieve a comprehensive evaluation.
DM unlearning methods to be benchmarked. In this work, we assess 9 most recently proposed MU methods for DMs, including ESD [23], FMN [28], UCE [24], CA [25], SalUn [27], SEOT [30], SPM [26], EDiff [31], and SHS [32]. We remark that SA [29], a recently proposed method, is excluded from our evaluation due to its excessive computational resource requirements and time consumption. Unless specified otherwise, SD v1.5 is the model used for unlearning. Detailed training settings for each method can be found in Appx. B.
4 Experiment Results
In this section, our benchmarking results are divided into two main parts. In Sec. 4.1, we comprehensively evaluate the existing DM unlearning methods on the established tasks of style and object unlearning. Extensive studies following the proposed evaluation pipeline (Fig. 5) reveal that focusing solely on unlearning effectiveness can lead to a biased perspective for DM unlearning if retainability is not concurrently assessed (Tab. 2 & Fig. A2). We also show that preserving CRA (cross-domain retainability) is more challenging than IRA (in-domain retainability), highlighting a gap in the current literature (Fig. 6). Furthermore, we find that different DM unlearning methods exhibit distinct unlearning mechanisms by examining their unlearning directions (Fig. 7). In Sec. 4.2, we introduce more challenging unlearning scenarios. First, we assess the robustness of current DM unlearning methods against adversarial prompts crafted based on (Fig. 8). Additionally, we examine the performance when facing unlearning targets with finer granularity, formed by style-object concept combinations (Tab. 3). Further, we leverage UnlearnCanvas to provide insights into DM unlearning in a sequential unlearning fashion (Tab. A5).
4.1 Benchmarking Current DM Unlearning Methods for Style and Object Unlearning
Overall performance of DM unlearning.
In Tab. 2, we provide an overview of the performance of existing unlearning methods, using the evaluation metrics associated with UnlearnCanvas.
@@ 4.1 Benchmarking Current DM Unlearning Methods for Style and Object Unlearning
Table 2: Performance overview of different DM unlearning methods evaluated on UnlearnCanvas. The performance metrics include UA (unlearning accuracy), IRA, CRA, and FID. The symbols $\uparrow$ or $\downarrow$ indicate whether a higher or lower value represents better performance. Results are averaged over all the style and object unlearning cases. The best performance for each metric is highlighted in green, while significantly underperforming results are marked in red, indicating areas needing improvement for existing DM unlearning methods.
@@
ESD [23] ${98.58}\%$ $80.97\%$ $93.96\%$ $92.15\%$ ${55.78\%}$ ${44.23\%}$ $65.55$ 6163 17.8 4.3
FMN [28] $88.48\%$ ${56.77\%}$ ${46.60\%}$ ${45.64\%}$ $90.63\%$ $73.46\%$ 131.37131.37{131.37}131.37 $350$ 17.9 4.2
UCE [24] ${98.40}\%$ ${60.22\%}$ ${47.71\%}$ ${94.31}\%$ ${39.35\%}$ ${34.67\%}$ 182.01182.01{182.01}182.01 434 5.15.1{5.1}5.1 1.71.7{1.7}1.7
CA [25] ${60.82\%}$ ${96.01}\%$ $92.70\%$ ${46.67\%}$ $90.11\%$ $81.97\%$ 54.2154.21{54.21}54.21 734 10.1 4.2
SalUn [27] $86.26\%$ $90.39\%$ $95.08\%$ $86.91\%$ ${96.35}\%$ ${99.59}\%$ $61.05$ 667 30.8 4.0
SEOT [30] ${56.90\%}$ $94.68\%$ $84.31\%$ ${23.25\%}$ $95.57\%$ $82.71\%$ 62.38 9595{95}95 7.34 0.0
SPM [26] ${60.94\%}$ $92.39\%$ $84.33\%$ $71.25\%$ $90.79\%$ $81.65\%$ 59.79 29700 6.9 0.0
EDiff [31] $92.42\%$ ${73.91\%}$ ${98.93}\%$ $86.67\%$ $94.03\%$ ${48.48\%}$ 81.42 1567 27.8 4.0
SHS [32] $95.84\%$ $80.42\%$ ${43.27\%}$ $80.73\%$ $81.15\%$ ${67.99\%}$ 119.34 1223 31.2 4.0
@@
First, as seen in Tab. 2, retainability is essential for a comprehensive assessment of DM unlearning. For example, in the scenario of style unlearning, ESD and UCE achieve similar UA (both around $98\%$), but their retainability (IRA and CRA) differs significantly, with one over $80\%$ and the other around $60\%$. Therefore, relying solely on UA can provide a skewed view of the performance of DM unlearning. Second, we observe that CRA is harder to retain than IRA. For example, UCE exhibits a significant gap between its IRA and CRA, both in style ($60.22\%$ vs. $47.71\%$) and object ($39.35\%$ vs. $34.67\%$) unlearning scenarios. Similar patterns are observed with other methods like FMN, CA, SEOT, SPM, and SHS. This pattern suggests that while MU methods are somewhat effective at preserving concepts within the same domain, they face greater challenges in maintaining performance on unlearning target-unrelated concepts across domains. This issue has been overlooked in previous studies due to the absence of a multi-label dataset like UnlearnCanvas for systematic unlearning analyses. Third, we find that a single unlearning method can perform differently across various domains, and no single method excels in all aspects. For example, FMN achieves a UA of 88.48% in style unlearning but a much lower UA of 45.64% in object unlearning. A similar phenomenon can be observed with the unlearning methods SEOT, SPM, and SHS, which exhibit significantly large performance gaps between style and object unlearning. Furthermore, each method exhibits unique strengths and weaknesses. For example, while ESD typically shows improvement in UA, it may lag in terms of IRA and CRA, particularly in the case of object unlearning. Some methods, such as SEOT and SPM, achieve extremely high storage efficiency, requiring no additional storage, but suffer from inferior UA (below 80%). These observations underscore the challenges of DM unlearning and highlight the need for further advancements in the field.
Figure 6: Left: Heatmap visualization of the unlearning accuracy and retainability of ESD on UnlearnCanvas. The $x$-axis shows the tested concepts for image generation using the unlearned model, while the $y$-axis indicates the unlearning target. Concept types are distinguished by color: styles in blue and objects in orange. The figure is divided into regions representing corresponding evaluation metrics and unlearning scopes ($A$ for UA, $B$ for IRA, $C$ for CRA; ‘1’ for style unlearning, ‘2’ for object unlearning). Higher values in lighter colors denote better performance. The first row serves as a reference for comparison before unlearning. Zooming into the figure is recommended for detailed observation. Right: Representative cases illustrating each region with images generated before and after unlearning a specific concept.
A closer look into retainability: A case study on ESD.
We next perform an in-depth analysis of DMs’ retainability after unlearning, focusing on the most popular ESD method [23]. In Fig. 6 (left), we present a heatmap illustrating the per-style/object unlearning accuracy (diagonal) and retain accuracy (off-diagonal) for ESD. The heatmap segments are labeled to indicate different regions of interest (ROIs) in style and object unlearning related to our evaluation metrics: A1/A2 for UA, B1/B2 for IRA, and C1/C2 for CRA. Key observations are highlighted below.
First, we observe distinct behaviors of ESD in unlearning styles compared to objects. For example, ROI C2 appears darker than C1, suggesting a reduced ability to retain styles during object unlearning. Similar patterns are seen between B2 and B1. For a clearer explanation, Fig. 6 (right) provides image generation examples of the DM before and after unlearning within different ROIs. Second, We observe that style/object unlearning is relatively easier compared to retaining the generation performance of unlearned DMs conditioned on unlearning-unrelated prompts. This is evident in the contrast between ROIs A1/A2 and ROIs B1/B2 or C1/C2. One exception is the case of unlearning the object ‘Sea’, which exhibits relatively low UA in A2. This may be attributed to post-unlearning image generation still resembling sea waves, as shown by the image examples in Fig. 6 (right) for A2. Third, we observe an inherent trade-off between unlearning effectiveness and retainability. For example, while SalUn exhibits more consistent performance across different scenarios than ESD, it does not achieve the same level of UA, as indicated by the lighter diagonal values in Fig. A3. For the heatmap performance of other unlearning methods, please refer to Appx. D.
@@
Understanding unlearning method’s behavior via unlearning directions.